OpenStudy (anonymous):

find the following: (k-g)(5) if g(x)= square root of x-3 and k(x)=x2+5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

subtract function and then plug in number

OpenStudy (anonymous):

Is your question: 5*( sqrt{x-3} - (x²+5))=?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

really, are you sure it is not F[5]=Sqrt[x-3]-x^2+5

OpenStudy (anonymous):

Either way, it contains complex numbers so am not even gonna bother.

OpenStudy (anonymous):

yes this is the problem. (k-g)(5) if k(x)= x^2+5 g(x)=sqrt{x-3}

OpenStudy (llort):

You substitute. if you know what g and k equal, or at least have a formula, you can plug it in to the original equation.